A series circuit of a resistor R = 1.40×101 Ω, an inductor L = 9.00×10-2 H, and a battery of emf = 6 Volts has a time constant tL. The circuit is closed at time t = 0. What is the time constant of this circuit? What is the value of the current in the circuit when t = 3.21E-3 s? When t = 1.29E-2 s?When t = 2.57E-2 s?
Answer #1

Time constant τ = L/R = 0.09/14 = 6.42e-3

a)I = V/R [1- exp (-t/τ)]

I = (6/14) (exp (1-3.21e-3/ 6.42e-3)

I = 0.168 A

b)  I = V/R [1- exp (-t/τ)]

I = (6/14) (exp (1-1.29e-2/ 6.42e-3)

I = 0.37 A

c) I = V/R [1- exp (-t/τ)]

I = (6/14) (1-exp (-2.57e-2/ 6.42e-3)

I = 0.42 A